The unit digit of a two-digit number exceeds twice the tens digit by 1. Find the number if the sum of its digits is 10.
Answer by drk(1908)

u = 2t + 1
u + t = 10
by substitution, we get
2t + 1 + t = 10
3t + 1 = 10
3t = 9
t = 3
u = 7
---
the number is 37